Suppose the risk-free rate is 4.3 percent and the market portfolio has an expected return of 11 percent. The market portfolio has a variance of .0392. Portfolio Z has a correlation coefficient with the market of .29 and a variance of .3295 According to the capital asset pricing model, what is the expected return on Portfolio Z? (Do not round intermediate calculations and enter your answer as a percent rounded to 2 decimal places, e.g., 32.16
Answer #1

Bela- covasiance (l,m) var.(m) Cosselation = covariance lim) of coefficient si sm - 0.gg = covariance (im) 50.3295 10.0392 =Ry - Rf + (Rm. Rf x Bela = 4.3+(11-4.3)%0.840780 = 9.933226 - 19.93%